OpenStudy (christos):

Find the area A of the sector of a circle of radius r=4 inches and central angle Θ=5 radians

OpenStudy (tkhunny):

Area of Whole Circle: \(\pi 4^{2}\) Central Angle of Whole Circle: \(2\pi\)

OpenStudy (christos):

any formula please... ?

OpenStudy (christos):

@tkhunny

OpenStudy (tkhunny):

I'm hoping you will think about it and come up with the answer. How much of the circle are you selecting when you select an angle of 5 Radians? Kowing that the Whole Circle is \(2\pi\) Radians, how does that compare to 5 Radians?